# A Lanczos-Based Algorithmic Approach for Spike Detection in Large Sample Covariance Matrices

### v1

@@ Acknowledgements
This material is based upon work supported by NSF DMS-2306438 (TT), DMS-2306439 (XD). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author and do not necessarily reflect the views of the National Science Foundation. Parts of this work were completed on Hyak, UW’s high performance computing cluster.
@@ A.2. Matrix Identities
Suppose $p,q$ are nonnegative integers such that $p+q>0$, and suppose $A,B,C,D$ are respectively $p\times p$, $p\times q$, $q\times p$, $q\times q$ matrices of complex numbers. Let
@@ Appendix B Proofs of some techinal lemmas
[…]
@@ Appendix C Pole computation
We present an approach for computing the poles of the Stieltjes transform associated with perturbations of Toeplitz Jacobi operators through connection coefficients, as described in [50]. An alternative method to approximate these poles, by truncating the semi-infinite matrix, is also presented here. The first method, that of Olver & Webb, gives guaranteed accuracy in exact arithmetic. The second offers weaker guarantees but offers a simpler implementation. Other techniques, such as those in [7], may also be useful in this context.
@@ C.1. Finite rank perturbation of Jacobi operators and connection coefficients
where ${\left\langle\cdot,\cdot\right\rangle}_{\mu}$ is the inner product for $L^{2}(\mu)$. The significance of this matrix lies in its ability to connect the spectral properties of $\widetilde{\mathcal{J}}$ with those of $\mathcal{J}$, allowing the spectrum of $\widetilde{\mathcal{J}}$ and its spectral measure to be fully determined. It is worth noting that the entries of $C$ follow a well-defined recurrence relation. For further details, see [44], [50, Lemma 3.2] and [51].
Lemma C.1 ([50], Lemma 3.2).
@@
Due to the structured form of $\widetilde{\mathcal{J}}$ in (C.3), (C.4), the connection coefficients matrix also inherits a specific structure. In particular, it decomposes as [50],
@@
where $C_{\mathrm{Toe}}$ is a Toeplitz matrix with bandwidth $12n-$ 1, and $C_{\mathrm{fin}}$ has nonzero entries only within the $(n-1)\times(n-2)$ principal submatrix. The next theorem provides a precise description and demonstrates that the entries of $C$ can be computed in $\mathrm{O}(n^{2})$ operations.
Theorem C.2 ([50], Theorem 4.8).
@@
(C.8) $c_{i,j}=c_{i-1,j-1}\text{ for all }i,j>0\text{ such that }i+j\geq 2n,\quad\text{and}\quad c_{0,j}=0\text{ for all }j\geq 2n.$
The connection coefficient matrix $C_{\widetilde{\mathcal{J}}\to\mathcal{J}}$ provides a way to find the poles of the Stieltjes transform associated with $\widetilde{\mathcal{J}}$ and to express its spectral measure in terms of that of $\mathcal{J}$. For a more detailed discussion, refer to [50, Theorems 4.12, 4.14 and Remark 4.15].
@@
can be used to overcome these issues. The Joukowski map is a conformal map from $\mathbb{D}=\{z\in\mathbb{C}:|z|<1\}$ to $\mathbb{C}\setminus[-1,1]$ that sends the unit circle to two copies of the interval $[-1,1]$.
Theorem C.4 ([50], Theorems 4.21 and 4.22).
@@
where $\{z_{i}\}_{i=1}^{r}$ are the roots of $c$ that lie in $\mathbb{D}$, which are all real and simple, and $M_{[\gamma_{-},\gamma_{+}]}$ is the affine map from $[-1,1]$ to $[\gamma_{-},\gamma_{+}]$ with $\gamma_{\pm}$ defined in (C.2). Moreover, the spectral measure of $\widetilde{\mathcal{J}}$ is
@@
The key ideas of this section are summarized in Algorithm PE.1. As established in [50, Theorem 6.8], this pole detection algorithm determines the exact number and location of the poles with arbitrary accuracy in a finite number of operations.
@@ C.2. Finite section of Jacobi operators
[…]
where $\gamma_{\pm}=\alpha\pm 2\beta$ and $\lambda_{1}\geq\dots\geq\lambda_{r}>\gamma_{+}$, and further assume that it sastisfies Assumption 1. The finite-section method [23, 24] is a common approach to estimating the spectrum of a linear self-adjoint operator. It approximates the spectrum by truncating the associated semi-infinite matrix to an $N\times N$ principal submatrix and using its eigenvalues for sufficiently large $N$. Since the poles of the Stieltjes transform associated with $\mathcal{J}$ correspond to the discrete eigenvalues $\{\lambda_{i}\}_{i=1}^{r}$, the finite-section method, as described in Algorithm PE.2 can be applied to estimate the number and locations of these poles.
Although this method is straightforward, it often has a significant drawback, known as spectral pollution. This occurs when the eigenvalues of the truncated operators (or a subsequence) converge to values that are not part of the true spectrum of the original operator. However, under the current assumptions, it can be shown that the finite-section method does not suffer from this issue and accurately captures the correct number of discrete eigenvalues when $N$ is sufficiently large. This can be established, for example, using the asymptotics of the orthogonal polynomials associated with $\mathcal{J}$, as outlined in [15]. The convergence of individual eigenvalues occurs at an exponential rate as $N$ increases. The rate does potentially degenerate as gaps in the spectrum of $\mathcal{J}$ close. A more detailed analysis of this observation is left for future work.

### v2

@@ Acknowledgements
This material is based upon work supported by NSF DMS-2306438 (TT), DMS-2306439 (XD). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author and do not necessarily reflect the views of the National Science Foundation.
@@ A.2. Matrix Identities
Suppose $p,q$ are nonnegative integers such that $p+q>0$, and suppose $A,B,C,D$ are respectively $p\times p$, $p\times q$, $\times p, q×q$$\times q$ matrices of complex numbers. Let
@@ Appendix B Proofs of some techinal lemmas
This establishes (4.26) by observing that $\widehat{\mathbf{u}}=\sum_{i=r+1}^{N}{u}_{i}\mathbf{v}_{i}$ and $\widehat{\mathbf{v}}=\sum_{i=r+1}^{N}{v}_{i}\mathbf{v}_{i}$.
@@ Appendix C Pole computation
We present an approach for computing the poles of the Stieltjes transform associated with perturbations of Toeplitz Jacobi operators through connection coefficients, as described in [Webb2021]. An alternative method to approximate these poles, by truncating the semi-infinite matrix, is also presented here. The first method, that of Olver & Webb, gives guaranteed accuracy in exact arithmetic. The second offers weaker guarantees but offers a simpler implementation. Other techniques, such as those in [bini_computing_2023], may also be useful in this context.
@@ C.1. Finite rank perturbation of Jacobi operators and connection coefficients
where ${\left\langle\cdot,\cdot\right\rangle}_{\mu}$ is the inner product for $L^{2}(\mu)$. The significance of this matrix lies in its ability to connect the spectral properties of $\widetilde{\mathcal{J}}$ with those of $\mathcal{J}$, allowing the spectrum of $\widetilde{\mathcal{J}}$ and its spectral measure to be fully determined. It is worth noting that the entries of $C$ follow a well-defined recurrence relation. For further details, see [sack_algorithm_1971], [Webb2021, Lemma 3.2] and [wheeler_modified_1974].
Lemma C.1 ([Webb2021], Lemma 3.2).
@@
Due to the structured form of $\widetilde{\mathcal{J}}$ in (C.3), (C.4), the connection coefficients matrix also inherits a specific structure. In particular, it decomposes as [Webb2021],
@@
where $C_{\mathrm{Toe}}$ is a Toeplitz matrix with bandwidth 2n12n-1, and $C_{\mathrm{fin}}$ has nonzero entries only within the $(n-1)\times(n-2)$ principal submatrix. The next theorem provides a precise description and demonstrates that the entries of $C$ can be computed in $\mathrm{O}(n^{2})$ operations.
Theorem C.2 ([Webb2021], Theorem 4.8).
@@
[…]
The connection coefficient matrix $C_{\widetilde{\mathcal{J}}\to\mathcal{J}}$ provides a way to find the poles of the Stieltjes transform associated with $\widetilde{\mathcal{J}}$ and to express its spectral measure in terms of that of $\mathcal{J}$. For a more detailed discussion, refer to [Webb2021, Theorems 4.12, 4.14 and Remark 4.15].
@@
can be used to overcome these issues. The Joukowski map is a conformal map from $\mathbb{D}=\{z\in\mathbb{C}:|z|<1\}$ to $\mathbb{C}\setminus[-1,1]$ that sends the unit circle to two copies of the interval [1,1][-1,1].
Theorem C.4 ([Webb2021], Theorems 4.21 and 4.22).
@@
where $\{z_{i}\}_{i=1}^{r}$ are the roots of $c$ that lie in $\mathbb{D}$, which are all real and simple, and $M_{[\gamma_{-},\gamma_{+}]}$ is the affine map from [1,1][-1,1] to $\gamma_{-},\gamma_{+}]$ with $\gamma_{\pm}$ defined in (C.2). Moreover, the spectral measure of $\widetilde{\mathcal{J}}$ is
@@
The key ideas of this section are summarized in Algorithm PE.1. As established in [Webb2021, Theorem 6.8], this pole detection algorithm determines the exact number and location of the poles with arbitrary accuracy in a finite number of operations.
@@ C.2. Finite section of Jacobi operators
(C.13) $\mu(\mathrm{d}\lambda)=h(\lambda)\sqrt{\lambda-\gamma_{-}}\sqrt{\gamma_{+}-\lambda}\,\mathds{1}_{[\gamma_{-},\gamma_{+}]}(\lambda)+\sum_{i=1}^{r}w_{i}\delta_{\lambda_{i}},$
where $\gamma_{\pm}=\alpha\pm 2\beta$ and $\lambda_{1}\geq\dots\geq\lambda_{r}>\gamma_{+}$, and further assume that it sastisfies Assumption 1. The finite-section method [gohberg_convolution_2005, hagen_c*-algebras_2001] is a common approach to estimating the spectrum of a linear self-adjoint operator. It approximates the spectrum by truncating the associated semi-infinite matrix to an $N\times N$ principal submatrix and using its eigenvalues for sufficiently large $N$. Since the poles of the Stieltjes transform associated with $\mathcal{J}$ correspond to the discrete eigenvalues $\{\lambda_{i}\}_{i=1}^{r}$, the finite-section method, as described in Algorithm PE.2 can be applied to estimate the number and locations of these poles.
Although this method is straightforward, it often has a significant drawback, known as spectral pollution. This occurs when the eigenvalues of the truncated operators (or a subsequence) converge to values that are not part of the true spectrum of the original operator. However, under the current assumptions, it can be shown that the finite-section method does not suffer from this issue and accurately captures the correct number of discrete eigenvalues when $N$ is sufficiently large. This can be established, for example, using the asymptotics of the orthogonal polynomials associated with $\mathcal{J}$, as outlined in [Ding2021b]. The convergence of individual eigenvalues occurs at an exponential rate as $N$ increases. The rate does potentially degenerate as gaps in the spectrum of $\mathcal{J}$ close. A more detailed analysis of this observation is left for future work.